Dr. James E. Guilkey

Schlumberger Technology Corporation 1935 S. Fremont Drive Salt Lake City, Utah 84104

(801) 635-7351 jguilkey@slb.com

June 17, 2010

Dr. Satya Atluri Computer Methods in Engineering and Sciences Tech Science Press

Dear Dr. Atluri:

This letter is intended to accompany the recently submitted manuscript “Enhanced Understanding of Particle Simulations Through Deformation-Based Visualization”, Paper No. CMES201006171679.

This paper represents original work by the authors, all of whom have signed below. This work has not been published elsewhere, nor is it being submitted elsewhere. This letter and the accompanying signatures convey permission to publish this work in Computer Methods in Engineering and Sciences.

Contact information for each author, and their signature, follows.

A.N.M. Imroz Choudhury  Dr. Michael D. Steffen  Dr. Steven G. Parker
SCI Institute, Rm 3750  SCI Institute, Rm 3750  NVIDIA Corporation

72 S Central Campus Dr. 72 S Central Campus Dr. 2150 S 1300 E, Suite 500 Salt Lake City, UT 84112 Salt Lake City, UT 84112 Salt Lake City, UT 84106

(801) 585-1874  (801) 581-8596  (801) 674-7946
roni@cs.utah.edu  msteffen@cs.utah.edu  sparker@nvidia.com

We look forward to timely and insightful reviews of our manuscript.

Sincerely,

James E. Guilkey